What happens during an exothermic reaction?

The surroundings lose energy to the system.
The system loses energy to its surroundings.
No energy is transferred between the system and the surroundings.
Energy is lost from the system and the surroundings.
Which of the following correctly lists the energies from most to least?

nuclear, chemical, physical
physical, chemical, nuclear
physical, nuclear, chemical
chemical, physical, nuclear
How much energy is required to raise the temperature of 10.5 g of iron metal (c = 0.444 J/(g•°C)) from 25.2 °C to 30.5 °C?

4.66 J
24.7 J
232 J
2.35 J
1
Expert's answer
2020-05-26T14:21:37-0400

During an exothermic reaction the system loses energy to its surroundings.


Energies from most to least:

nuclear, chemical, physical


Energy required to raise the temperature of 10.5 g of iron metal (c = 0.444 J/(g•°C)) from 25.2 °C to 30.5 °C: Q = c*m*(T2-T1) = 0.444*10.5*(30.5-25.2) =

24.7 J